A mule and donkey were going to market laden with wheat .The mule said "If you give me one measure, I should carry twice as much as you , but if I give you one . We should have equal burdens "what was each of the burdens"
Answer by RAY100(1637)

(M+1)=2D
M=D+1
subst
D+1+1 = 2D
simplify by subst D both sides
2=D
then M=2+1 = 3
check
3+1 =2 *2 =4 ok
3= 2+1 ok